OpenStudy (is3535):

For 1 month, two doctors in a medical office counted the number of patients they saw each day. The mean and MAD for each doctor were calculated and recorded in a table. What conclusion can be drawn, based upon the value of the means-to-MAD ratio? Doctor Mean MAD Hartmann 15.3 0.9 Kim 18.9 1.1

OpenStudy (is3535):

A. Dr. Hartmann saw more patients per day than Dr. Kim did. B. Dr. Kim saw more patients per day than Dr. Hartmann did. C. Dr. Kim and Dr. Hartmann saw about the same number of patients per day. D. No conclusions can be drawn.

OpenStudy (black_ninja123):

What do you think it is?

OpenStudy (is3535):

i think C

OpenStudy (black_ninja123):

I would say B because Dr.Kim has greater numbers than Dr.Hartmann

OpenStudy (is3535):

ok I am not 100% sure, but I know how to do the problem, i think lol so mean "to" mad means divide your mean (15.3) by your MAD (0.9) etc etc so Dr. Hartman seen 15.3/0.9 = 17 patients Dr. Kim seen 18.9 / 1.1= 17.18 repeated patients. so they both aw about 17 patients, or about the same amount each so what should the answer be C

OpenStudy (black_ninja123):

I guess you should go with C

OpenStudy (is3535):

@Vocaloid

OpenStudy (is3535):

@iGreen

OpenStudy (is3535):

@Thesmarterone

OpenStudy (anonymous):

Lol you just copied and paste I am not 100% sure, but I know how to do the problem, i think lol so mean "to" mad means divide your mean (15.3) by your MAD (0.9) etc etc so Dr. Hartman seen 15.3/0.9 = 17 patients Dr. Kim seen 18.9 / 1.1= 17.18 repeated patients. so they both aw about 17 patients, or about the same amount each so what should the answer be? :) But changed it from someone else

OpenStudy (is3535):

ok I am not 100% sure, but I know how to do the problem, i think lol so mean "to" mad means divide your mean (15.3) by your MAD (0.9) etc etc so Dr. Hartman seen 15.3/0.9 = 17 patients Dr. Kim seen 18.9 / 1.1= 17.18 repeated patients. so they both aw about 17 patients, or about the same amount each so what should the answer be C by I.L